### v1

# Are Large Language Models Rational Investors?

Abstract
Large Language Models (LLMs) are progressively being adopted in financial analysis to harness their extensive knowledge base for interpreting complex market data and trends. However, their application in the financial domain is challenged by intrinsic biases (i.e., risk-preference bias) and a superficial grasp of market intricacies, underscoring the need for a thorough assessment of their financial insight. This study introduces a novel framework, Financial Bias Indicators (FBI), to critically evaluate the financial rationality of LLMs, focusing on their ability to discern and navigate the subtleties of financial information and to identify any irrational biases that might skew market analysis.
Our research adopts an innovative methodology to measure financial rationality, integrating principles of behavioral finance to scrutinize the biases and decision-making patterns of LLMs. We conduct a comprehensive evaluation of 19 leading LLMs, considering factors such as model scale, training datasets, input strategies, etc. The findings reveal varying degrees of financial irrationality among the models, influenced by their design and training. Models trained specifically on financial datasets might exhibit greater irrationality, and it’s possible that even larger financial language models (FinLLMs) could display more biases than smaller, more generalized models. This outcomes provide profound insights into how these elements affect the financial rationality of LLMs, indicating that targeted training and structured input methods could improve model performance. This work enriches our understanding of LLMs’ strengths and weaknesses in financial applications, laying the groundwork for the development of more dependable and rational financial analysis tools.
[…]
Are Large Language Models Rational Investors?
Yuhang Zhou††thanks: Email: yuhangzhou22@m.fudan.edu.cn Yuchen Ni Xiang Liu Jian Zhang Sen Liu Guangnan Ye††thanks: Corresponding Author. Email: yegn@fudan.edu.cn Hongfeng Chai Institute of Fintech, Fudan University School of Computer Science, Fudan University School of Electronics and Information Engineering, Tongji University Tandon School of Engineering, New York University DataGrand Inc.
1 Introduction
Recent advancements in LLMs, such as GPT-4 OpenAI (2023) and LLaMA-2 Touvron et al. (2023), have shown their prowess across a spectrum of natural language processing (NLP) tasks Zhao et al. (2023), extending into specialized domains such as finance Zhang et al. (2023), law Cui et al. (2023), and healthcare Wang et al. (2023). Despite their versatility, these models grapple with inherent biases Gallegos et al. (2023), Rutinowski et al. (2023), encompassing gender, race, and socioeconomic disparities, which could compromise their reliability and entail significant consequences Jeoung et al. (2023). Efforts to mitigate such biases have led to the development of benchmark datasets such as StereoSet Nadeem et al. (2020) for stereotype identification and OpinionGPT Haller et al. (2023) for generating bias-neutral content, employing techniques such as LoRA Hu et al. (2021) for model fine-tuning, especially within the social sciences. As we delve into the application of these advanced models in specific sectors, it becomes imperative to address these biases to ensure fair and unbiased outcomes.
In the realm of financial LLMs (FinLLMs), research has predominantly concentrated on enhancing model performance through continued pre-training or fine-tuning, with evaluation metrics focused on NLP tasks and financial applications. BloombergGPT Wu et al. (2023), a pioneering FinLLM, was trained on a vast financial corpus using BLOOM Workshop et al. (2022), and assessed across a broad array of financial and general NLP tasks. The FinQA dataset Chen et al. (2021), comprising QA pairs derived from S&P 500 company reports, served as a notable example of domain-specific evaluation. Moreover, LLMs have been employed in financial analytics, acting as advisors by leveraging news Lopez-Lira and Tang (2023) and fundamental analyses Fatouros et al. (2024) for investment decisions across diverse securities Romanko et al. (2023). However, the efficacy of these models is contingent upon the models’ own rationality as participants in the market. A lack of rationality in large models could lead to misinterpretations and misapplications of market dynamics, adversely impacting not only the users of these models but also the broader economy. Figure 1 is an example of model irrationality, where the model gives inconsistent expectations for the same event from different subjects, resulting in different emotions.
Hence, it is imperative to examine the rationality of LLMs before their integration into financial advisory roles. Limited research has addressed the financial biases in pre-trained models, with methodologies such as probabilistic detection Chuang and Yang (2022) and consistency checks Yang et al. (2023) being explored. However, these approaches primarily focus on data processing and lack a comprehensive financial theory to substantiate their findings, rendering them inadequate for assessing LLM financial rationality. This underscores the critical need for a holistic framework to evaluate the rationality of LLMs in financial contexts.
To work towards this goal, our study conducts an examination of the financial rationality of LLMs based on the theory of behavioral finance Barberis and Thaler (2003). We believe that the enduring theory of behavioral finance, grounded in psychology and finance, can provide a more comprehensive perspective to support research findings. In the current scenario where it is challenging to quantify behavioral finance indicators, we propose the Financial Bias Indicators (FBI) framework, which provides a comprehensive assessment of financial rationality within LLMs from multiple perspectives.
Our framework specifically delineates the financial rationality of LLMs into two dimensions: Belief and Risk Preference. By leveraging theories on human cognitive biases, we assess LLMs’ rationality by observing their reactions to phenomena, e.g., limited attention bias, anchoring effect, and loss aversion, thereby gauging their rationality. Additionally, we delve into the models’ reasoning processes in financial contexts to distinguish whether observed biases arise from inherent capabilities or a lack of robustness. This approach enables us to discern the nuanced behaviors of LLMs, revealing a spectrum of financial irrationality across models of varying sizes. Notably, our research uniquely focuses on the biases of models within China’s A-share market, a domain scarcely explored in existing literature, thereby filling a critical gap in the understanding of LLM behavior in specific financial environments.
Our findings indicate that both larger and smaller Foundation Models111Foundation Models is an alias for Large Language Models, which is used interchangeably in this article. manifest degrees of financial irrationality, with larger models showing subtler biases and smaller ones more pronounced irrational behaviors across different companies, industries, and scenarios. This irrationality, potentially deepened by continuous pre-training or incremental fine-tuning with financial datasets, poses a risk of market anomalies when such models are applied in real-world financial scenarios. Despite the potential for mitigation through instructional methods or reasoning approaches such as Chain of Thought (COT) Wei et al. (2022), the inherent biases in LLMs, underscore the need for further research aimed at enhancing the models’ robustness, fairness, and rationality to safeguard financial market stability and protect personal assets.
Our key contributions to this field are summarized as follows:
We have found that LLMs often exhibit irrational phenomena in the field of financial quantification research. So we introduce the Financial Bias Indicators (FBI) framework, a novel approach for assessing the rationality of LLMs in finance. Grounded in behavioral finance, the FBI framework categorizes financial rationality into Belief Bias and Risk-preference Bias, offering a new perspective for evaluating LLM rationality in financial contexts. To our best knowledge, this is the first study on the rationality of LLMs in the financial domain, and also the first study to quantify behavioral finance indicators, which can help future researchers develop more reliable FinLLMs (section 3). We have published our code and data on GitHub222https://github.com/zhiqix/FBI.
The FBI framework conducts an extensive analysis of the top 19 Chinese and English oriented LLMs, examining the impact of model parameters, training data, and input formats on their financial rationality. This comprehensive review provides insights into the varying degrees of financial irrationality across different models, enhancing our understanding of their behavior in financial applications(section 4 and section 5).
Based on the FBI framework, we further investigate the root causes of financial irrationality in LLMs from various perspectives. Gaining a deeper understanding of these factors enables more effective evaluation and mitigation of financial irrationality in LLMs, contributing to the creation of more robust and dependable models. This supports the advancement of global financial development(section 6).
2 Background
Behavioral finance Barberis and Thaler (2003), an area at the confluence of psychology and economics, investigates how psychological influences and cognitive biases affect the financial behaviors of individuals and institutions and, by extension, financial markets. This field seeks to elucidate the underlying psychological factors that lead to various market phenomena, diverging from traditional finance theories that assume rational market participants. By dissecting individual and collective financial decision-making processes, behavioral finance aims to construct a realistic framework of market dynamics, incorporating cognitive errors and the limits of arbitrage. We adopt the classification proposed by Barberis and Thaler (2003), segmenting behavioral finance into Cognitive Bias and Limits to Arbitrage, to structure our exploration of these dimensions.
2.1 Cognitive Bias
Cognitive biases represent systematic departures from normative decision-making, influencing how investors form beliefs and assess risks. These biases are multifaceted, manifesting as erroneous beliefs or inconsistent risk preference. For instance, belief biases, such as attentional neglect, representativeness bias, anchoring effect, and overconfidence,skew investors’ expectations, while risk-preference biases, such as loss aversion and reference dependence, lead to irregularities in risk assessment and decision-making under uncertainty. A comprehensive taxonomy of these biases, alongside their definitions, is detailed in Appendix B.
2.2 Limits to Arbitrage
Contrary to the Efficient Markets Hypothesis (EMH), which posits that asset prices fully reflect all available information and that market participants behave rationally, behavioral finance identifies scenarios where irrationality prevails, leading to anomalies like market bubbles and systemic crises. These phenomena are often attributed to the collective impact of cognitive biases on investor expectations, which can cause significant deviations from asset fundamentals. This section delves into how such irrational behaviors challenge the EMH, introducing inefficiencies that arbitrage mechanisms cannot always correct.
3 FBI: A Framework for Assessing LLMs Financial Rationality
We propose the FBI framework illustrated in Figure 2, drawing upon behavioral finance principles and the integration of NLP technologies in the financial sector. This framework bifurcates into two segments to analyzing biases in LLMs: Belief Bias and Risk-preferences Bias.
3.1 Belief Bias
In the contemporary milieu, the perpetual influx of novel information necessitates ongoing revisions to our forecasts and anticipations, shaping our beliefs. Within the belief bias examination, this framework endeavors to authenticate four biases: Anchoring Effect, Limited Attention Bias, Representativeness Bias, and Overconfidence, employing a uniform dataset. It evaluates the capacity of various LLMs to sustain objective rationality in actual market dynamics.
To evaluate the rationality of LLMs in financial markets comprehensively, we analyze the models’ responses to newly emerging information. This involves categorizing messages into two main types: event news and interactions, especially considering the noise prevalent in online investor dialogues within the Chinese market.
We adopt a refined classification of financial events, inspired by Zhou Zhou et al. (2021), into four primary categories: Corporate Governance and Equity Changes (CGEC), Financial Reports and Earnings Expectations (FREE), Market Behavior and Announcements (MBA), and Negative Events and Risk Management (NERM). This leads to a set of specific event types , as elaborated in Appendix B.
From January 2023 to December 2023, we compiled a corpus of news pieces, with a focus on including a subset that contains emotionally charged articles. This inclusion aims to enrich the LLM training dataset with event news that embodies emotional dimensions. The selected subset comprises 9 articles with positive emotions, 9 with negative emotions, and 6 with mixed emotions, resulting in a total of articles earmarked for detailed analysis.
Furthermore, we curated a set of 10 emotionally neutral question-response interaction pairs , to study LLMs’ comprehension in a controlled setting.
For each selected piece of news and interaction , we abstracted specific numerical details to proportional figures to standardize the data across varying scales of company market capitalizations.
We then sampled 600 companies from the Chinese A-share market, excluding those marked for delisting (ST), distributed equally across three tiers based on market capitalization: top, middle, and bottom, with companies from each tier. This selection was anonymized by replacing company and industry names to test the belief bias in LLMs, with a detailed classification by the first-level industry provided in Appendix C.
3.1.1 Analysis Logic
Our analysis framework is designed to unravel various belief biases potentially embedded within LLMs, using a structured approach to evaluate the models’ reactions to financial news across varying contexts.
Anchoring Effects
To detect anchoring effects, we examine whether LLMs display differing interpretations of identical events when presented with varying company contexts, indicative of bias ingrained from historical data. This involves modifying the subject company in each news item and assessing the variability in LLM evaluations using Analysis of Variance (ANOVA) St et al. (1989), denoted as , where tests the variance in scores across companies for a given news item . The specific prompt structure for LLMs is detailed in Appendix F.
Representativeness Bias
We explore the representativeness bias by correlating LLM output scores with company size and industry sector. The size correlation is quantified using the Spearman correlation coefficient , where represents the LLM scores and the market capitalizations of companies . Industry correlation analysis utilizes ANOVA, expressed as , alongside box plots to compare score distributions across different industries , representing the first-level industry classification of companies.
Overconfidence
To assess overconfidence, we observe score volatility for identical events with varied subjects among models undergoing continuous pre-training or fine-tuning on financial datasets. This is measured by the standard deviation of scores for the same event applied to different company contexts.
Limited Attention Bias
For detecting limited attention bias, we adopt a Comparative Output Testing (COT)-like methodology, allowing models to engage in a deliberative thought process before issuing scores. This process aims to discern if a more reflective approach versus a rapid response leads to more consistent scoring, analyzed through a variance comparison .
Score Instability Analysis
To investigate the roots of score instability, be it due to inadequate reasoning or compromised rationality, we utilize Bertopic Grootendorst (2022) for topic word extraction within the reasoning texts provided by the models. By clustering reasoning texts and extracting keywords, we analyze score discrepancies across different clusters, denoted by , to identify if certain thematic focuses lead to inconsistent evaluations.
3.2 Risk-preference Bias
Asset returns are inherently uncertain, and investor behavior within financial markets is influenced by their risk-return preferences. The exploration of Risk-preference Bias within our framework seeks to elucidate situational dependence bias, loss aversion, and the framing effect across diverse decision-making contexts. It evaluates the consistency of LLMs’ risk preferences amidst varying decision-making scenarios.
3.2.1 Data Design
In crafting scenarios that mirror real-life decision-making, we assembled a collection of 40 diverse situations spanning career planning, agricultural management, property transactions, and sporting events. Each scenario offers three decision alternatives , where , articulated in terms of potential gains or losses. The presentation order of these alternatives is randomized to minimize option bias in LLMs Zheng et al. (2023).
The construction of these alternatives adheres to the expected utility theory, encapsulated by the equation:
where is the utility function, the outcome under state , and the probability of .
Following the theoretical framework Simon et al. (1994), the utility function’s concavity mirrors risk preferences, with convexity indicating a different disposition. Considering the practical difficulties in direct concavity assessment, we employ the second-order Taylor expansion for utility approximation:
enabling us to modulate risk preferences through variance adjustment. This approach is detailed in Appendix E.
3.2.2 Analysis Logic
Recognizing decision-making as a process shaped by previous experiences and contextual factors, we delve into the situational dependence bias by examining if LLMs exhibit variable risk preferences across different scenarios . This involves prompting LLMs to choose among gain-framed scenarios and analyzing the pattern of choices across various contexts to discern any situational bias.
In the context of loss aversion, we scrutinize LLM responses within loss-framed scenarios , aiming to uncover any predominant risk-averse or risk-loving tendencies.
Furthermore, to investigate the framing effect, we translate all scenarios from Chinese to English, monitoring shifts in LLM preferences as a function of linguistic framing. This analysis seeks to identify if LLM outputs are influenced by the manner in which scenarios and choices are linguistically constructed, potentially revealing a bias towards the framing of options.
4 Models
In our evaluation, we select BGE333https://huggingface.co/BAAI/bge-large-zh-v1.5 as the Chinese embedding model and test 19 models, including English-oriented, Chinese-oriented, and FinLLMs, such as GPT4 OpenAI (2023), LLaMA Touvron et al. (2023), GLM4 Zeng et al. (2022), Qwen Bai et al. (2023),Xuanyuan Zhang et al. (2023), etc. The detailed settings refer to Appendix D.
5 Results
5.1 Belief Bias
The evaluation of Belief Bias is principally conducted through the examination of event news and interactions. This analysis reveals a widespread Anchoring Effect across the majority of Large Language Models (LLMs) when the subjects of events and interactions are modified, with slight variations observed among different models. The average variance index, detailed in Figure 3, sheds light on the rationality levels of LLMs with respect to representativeness bias. Specifically, LLMs with a focus on the Chinese language, such as the GLM and Qwen series, exhibit commendable financial rationality, whereas the Xuanyuan and Baichuan series are more susceptible to irrational behavior.
In terms of Overconfidence, violin plots, presented in subsection G.1, illustrate the score distributions of various texts across all models. A notable disparity is observed in the models’ responses to composite texts of positive and negative emotional content. As per Table 4, the GPT and InternLM models display a marked optimism, in contrast to the pronounced pessimism of the Qwen and GLM series. Furthermore, Figure 4 highlights that models trained on financial corpora experience a heightened score variability compared to their base counterparts.
In terms of Representativeness Bias, the less rational LLMs demonstrated a correlation coefficient between model scores and market capitalization below 10%, indicating a minimal correlation. However, certain LLMs showed clear biases towards specific industries, as documented in subsection G.1. For example, the FinQwen model consistently allocated lower scores to the Steel and Banking sectors. A comprehensive analysis reveals that the Media, Steel, Banking, and Non-Banking Finance sectors frequently occupy the extreme ends of the scoring spectrum across different models, whereas the Computer Science and Automobile sectors generally maintain a middle ground, exhibiting relative stability.
For the limited Attention Bias, the application of the COT methodology, aimed at slowing down the cognitive process of models, resulted in a minor reduction in rationality among the more logical models but significantly diminished biases in the less rational ones, as shown in subsection G.2. The clustering and keyword analysis indicate that this intentional cognitive strategy improves the logical consistency and financial acumen of the model outputs. A comparison of the reasoning keywords for the top-performing GLM-4 model and the underperforming Baichuan2-7B model, illustrated in Figure 39 and Figure 39 respectively, reveals a more coherent logic in GLM’s reasoning compared to Baichuan2-7B. Moreover, the clustering of the models’ inferential outputs reveals substantial variations in the ratings assigned to each category. This divergence underscores that the irrationality exhibited by models in financial contexts stems more from their inherent cognitive processes than from their computational capabilities.
5.2 Risk-preference Bias
The exploration of Risk-preference Bias entails the examination of LLM decisions across varied scenarios. The compiled results, particularly for gain-framed queries, are tabulated in Table 10. A predominant trend among most models is the exhibition of distinct risk preferences in disparate scenarios, indicative of a pronounced Situational Dependence Bias. Nonetheless, prefacing prompts with an instruction of the model’s risk-averse nature significantly attenuates this bias. In the context of loss-framed queries, some models exhibit a pronounced Loss Aversion Bias like GPT-4, as shown in Table 9. Moreover, refer to Table 8, the translation of all queries into English precipitated notable discrepancies between the models’ responses to Chinese and English versions, underscoring a pronounced Framing Effect.
6 Discussion
Our study’s contribution through the FBI framework extends to the identification and reduction of irrational tendencies in large-scale models within the finance sector. By employing FBI, researchers can acquire an in-depth comprehension of the biases and logical processes of LLMs in finance. Such insights are pivotal for pinpointing detrimental financial irrationalities and devising strategies to mitigate them. We delve into the influence of Model Size, Training Data, and Input Form modifications on the observed outcomes.
6.1 Model Size
The findings from subsection 5.1 illustrate a correlation between the enlargement of model parameters and the enhancement of financial rationality, particularly in relation to belief bias. This trend underscores the significance of model size in advancing NLP within financial quantitative analysis. Therefore, our analysis shows that the financial rationality of LLMs often increases with the increase in model size, which is consistent with the scaling law Kaplan et al. (2020).
6.2 Training Data
The FBI framework evaluated the financial rationality of both general-LLMs and FinLLMs. Insights from Figure 4 indicate that models trained on financial datasets are prone to heightened score variability and an increased inclination towards risk, thereby potentially exacerbating financial irrationality. Meanwhile, based on figures from Figure 12 to Figure 29, we observe that many models exhibit nearly opposite industry biases, such as ChatGLM2-6B and Qwen-7B, yet models from the same period often share similar biases. This suggests that such biases are likely influenced by temporal biases in the training data, in line with the concept of industry cycle rotation. Financial datasets typically encompass research reports, financial statements, corporate disclosures, and financial regulations. Previous studies Wang and Wu (2011) have highlighted the potential for embellishment and over-promotion within online financial content, such as research reports. While incorporating such data can bolster LLM capabilities in NLP tasks specific to finance, it also harbors the risk of embedding financial irrationality. The deployment of these models in financial quantification tasks could inadvertently precipitate financial disturbances, posing a critical concern that necessitates immediate attention.
6.3 Input Forms
The FBI framework employed three prompt approaches: Direct, Instructional, and COT. Experimental findings from section 5 suggest that the inclusion of instructional prompts or the application of the COT method can enhance the financial rationality of models. Specifically, the Instructional method enables models to adopt predefined risk preference attributes, thereby improving their adherence to directives. Implementing the COT method to emulate a "slow thinking" process effectively moderates the models’ susceptibility to belief bias.
7 Conclusion
Our research introduces the Financial Bias Indicators (FBI) framework as a pioneering approach to critically assess the financial rationality of Large Language Models (LLMs) within the complex domain of financial analysis. Through a rigorous examination of 19 leading LLMs, our study unveils significant variances in their financial rationality. The findings underscore the pivotal role of model architecture, training paradigms, and input forms in shaping the financial rationality of LLMs. By highlighting the nuances of financial rationality and the impact of inherent biases, this work not only enhances our understanding of LLMs’ capabilities and limitations in financial contexts but also charts a path forward for refining these sophisticated tools. The development of LLMs that are both more financially astute and unbiased is crucial for their reliable application in financial analysis, promising a future where AI-driven insights can be leveraged with greater confidence and precision in the financial sector.
Appendix A Cognitive Bias
For cognitive bias, we classified it into Belief Bias and Risk reference Bias based on previous research, and studied seven of these biases within the FBI framework.Refer to Table 1 for specific content.
| Cognitive Bias | Bias Type | Definition |
| Belief Bias | Limited Attention |
The brain has two systems when working: fast thinking and slow thinking. It uses intuition to deal with things quickly. |
| Situational dependence bias |
The effect of a stimulus depends largely on the context in which it occurs. |
| Loss aversion |
Sensitivity to losses exceeds gains of equal value. |
| Framing effect |
Different descriptions of an objectively identical problem lead to different decision-making judgments. |
| Risk-Preference Bias | Representativeness bias |
When making probability estimates, people tend to focus on certain representative features, ignoring environmental probabilities and sample size. |
| Anchoring effect |
Decision-making is often influenced by the first information received, like an anchor sinking to the bottom of the sea. |
| Overconfidence |
Belief that one’s knowledge is more accurate than the facts; one’s information is given more weight. |
Appendix B Event Type
We have sorted out the types of events that can affect a company’s stock price based on the regular patterns of the Chinese stock market, and finally sorted out four categories, totaling 16 types of events. The detailed content is shown in Table 2.
| Event Type | Subdivision Type | Definition |
| Corporate Governance and Equity Changes |
Major Asset Restructuring |
The process of recombining, adjusting, and allocating the distribution status of enterprise assets among the owners, controllers, and external economic entities. |
Equity Incentive |
By conditionally granting employees partial shareholder rights, a sense of ownership is fostered, forming a community of interests with the company. |
Increase or Decrease in Shareholder Holdings |
Changes in the shareholder holdings of company stocks. |
Buy-back |
The act of a listed company using cash or other means to repurchase its shares from the stock market. |
Circulation of Restricted Stock |
Restricted shares become freely tradable in the secondary market after the commitment period. |
| Financial Reports and Earnings Expectations |
Performance Report |
Regular preparation by each responsibility center to evaluate and assess performance, serving as the basis for future budget preparation. |
| Market Behavior and Announcements |
Private Placement |
Targeted issuance of bonds or stocks to a select group of senior institutional or individual investors. |
Transfer of Shares |
Listed companies transfer their provident fund to share capital in proportion or issue bonus shares accordingly. |
Stock Price Fluctuations |
Sudden large inflows and outflows of funds lead to increased volatility in stock prices. |
Business Dynamics |
Updates on enterprises and their surroundings, using major production and sales information to promote corporate brand and image. |
| Negative Events and Risk Management |
Dispute |
Disputes between companies or between companies and individuals. |
Investigation |
Filing an investigation signifies a basic determination of illegal facts, allowing for compulsory measures and official initiation of investigation procedures. |
Violation Penalties |
Punishments for enterprises violating regulations of regulatory bodies. |
Litigation and Arbitration |
Litigation and arbitration for contract disputes and other property rights disputes between enterprises. |
Security |
Enterprises providing guarantees for loans and other matters for other enterprises. |
Appendix C Company Profile
In order to avoid bias caused by market value impact, we did not choose funds from the CSI 300 or CSI 500. Instead, we summarized all listed companies in China. After removing ST type stocks, we selected the top, middle, and bottom 200 stocks based on market value, totaling 600 stocks. The industry distribution of stocks is shown in the Figure 5.
Appendix D Models
We have selected a total of 19 financial and generic LLMs oriented by Chinese and English, with specific details shown in Table 3.
| Model Name | Chinese-oriented | Model size | FinLLM | Deployment method |
| Baichuan2-7B | True | 7B | False |
local |
| Baichuan2-13B | True | 13B | False |
local |
| ChatGLM2-6B | True | 6B | False |
local |
| ChatGLM3-6B | True | 6B | False |
local |
| ChatGLM3-Turbo | True | 33B | False |
API |
| GLM-4 | True | Unknown | False |
API |
| InternLM2-7B | True | 7B | False |
local |
| InternLM2-20B | True | 20B | False |
local |
| LLaMA2-7B | False | 7B | False |
local |
| LLaMA2-13B | False | 13B | False |
local |
| Qwen-7B | True | 7B | False |
local |
| Qwen-14B | True | 14B | False |
local |
| FinQwen | True | 14B | True |
local |
| Qwen-72B | True | 72B | False |
local |
| Qwen-max | True | 72B | False |
API |
| Xuanyuan-13B | True | 13B | True |
local |
| Xuanyuan-70B | True | 70B | True |
local |
| GPT-3.5 | False | Unknown | False |
API |
| GPT-4 | False | Unknown | False |
API |
Appendix E Formula Proof
Invoking the fundamental principles of expected utility theory, we recognize that a utility function’s curvature reflects an individual’s risk preference. Specifically, a concave utility function () is indicative of risk aversion, while a convex utility function () signifies risk-seeking behavior. A linear utility function (), on the other hand, corresponds to risk neutrality.
The expected utility can be formally represented as:
Here, denotes the utility function, symbolizes the outcome under state , and is the probability of outcome occurring.
Furthermore, we articulate the variance of outcomes , , as the expected squared deviation from the expected value :
Applying the second-order Taylor expansion to the utility function around the expected value furnishes us with:
Imposing expectations on the approximated function, we derive the expected utility approximation:
Since , the middle term vanishes, simplifying our expression to:
Consequently, under the assertion of utility function concavity or convexity, the sign of the second derivative establishes the nature of risk preference. For a negative second derivative (), indicative of risk aversion, a smaller variance is required to enhance the expected utility. Conversely, for a positive second derivative (), characteristic of risk-seeking behavior, a larger variance is preferred. Risk-neutral individuals () show indifference to the variance level.
Through this analytical framework, we delineate how the variance of outcomes in conjunction with the utility function’s concavity or convexity guides the determination of an individual’s risk preference.
Appendix F Framework for Data Construction
This section delineates the structured approach employed in the study to formulate datasets incorporating event news, interactive elements, and risk preference inquiries. Each category of information is meticulously crafted using a distinct template, which is elucidated below.
F.1 Event News Template
The construction of the event news dataset leverages prompt engineering techniques to embed real-world events within a framework that facilitates evaluation, simulating the analytical capabilities of financial experts. The evaluation process involves the model assigning a score to each event based on its potential positive or negative impact on the financial landscape. Initially, the model is instructed to provide an immediate, intuitive score reflecting a ’fast thinking’ approach.
To augment the depth of analysis and ensure the robustness of the evaluation, the model is further tasked with adopting a ’slow thinking’ strategy. This entails a comprehensive articulation of the rationale behind the score, encouraging a deliberate and reasoned assessment. The detailed format of this template is illustrated in Figure 6, which guides the model in delivering both the quantitative score and the qualitative reasoning underpinning it.
F.2 Interactions
Using input methods similar to news events for rewriting, the specific template is shown in Figure 7.
F.3 Risk-Preference Questionnaire Template
The methodology for assessing risk preferences through structured questions is twofold, designed to discern the inherent risk orientation of the AI model under different conditions. Initially, the model is presented with a set of scenarios where it must select an option that best aligns with its assessed risk profile, simulating an introspective decision-making process. This setup aims to capture the model’s spontaneous risk preferences without external biases.
Subsequently, the experiment introduces a predefined constraint by explicitly characterizing the model as risk-averse within the instructions. This manipulation is intended to observe the adaptability of the model’s responses to altered risk parameters, thereby evaluating its capacity for contextual behavioral adjustment. The layout and content of these questions are encapsulated in the template depicted in Figure 8, which systematically guides the model through the decision-making process under varying risk conditions.
Appendix G Result
G.1 Analysis of Direct News Events
The examination of direct news events involves a detailed statistical analysis of the responses generated by various Large Language Models (LLMs) to specific news items. This analysis primarily focuses on the distribution of scores assigned by LLMs to each news event, encompassing key statistical measures such as the mean, variance, highest, and lowest scores. Such an approach is instrumental in assessing the consistency and rationality of LLMs’ interpretations of financial news.
To facilitate a comprehensive understanding of these scoring distributions, this section will present violin plots for each news event. Violin plots offer a more nuanced visualization compared to traditional box plots by showing the probability density of the data at different values. This graphical representation will thus provide insights into the spread and skewness of LLMs’ ratings across various news events, enabling a deeper analysis of their evaluative patterns and potential biases.
ıin 0,2,4,…,22
In the analysis of the initial five events, the focus is placed on news items that encompass both positive and negative performance reports, alongside fluctuations in stock prices. This diverse array of news content allows for a multifaceted examination of each Large Language Model’s (LLM’s) scoring tendencies. Notably, discrepancies in scoring preferences among different LLMs emerge when confronted with this spectrum of financial news.
A systematic statistical analysis is conducted on the scoring outcomes attributed to the positive and negative aspects of these events. This entails a detailed examination of how each LLM assesses the same news piece, shedding light on the variance in their interpretations and the potential implications of their biases. The findings from this analysis are meticulously compiled and presented in Table 4, offering a clear, quantified insight into the LLMs’ evaluative patterns across the selected news events.
| Model | Positive Times |
| GPT-4 | 5 |
| InternLM2-20B | 5 |
| LLaMA2-13B | 5 |
| Qwen-72B | 5 |
| FinQwen | 4 |
| InternLM2-7B | 4 |
| LLaMA2-7B | 4 |
| Qwen-max | 4 |
| Xuanyuan-13B | 4 |
| Baichuan2-13B | 3 |
| Baichuan2-7B | 3 |
| ChatGLM3-Turbo | 3 |
| GLM-4 | 3 |
| Xuanyuan-70B | 3 |
| ChatGLM2-6B | 2 |
| ChatGLM3-6B | 2 |
| Qwen-14B | 2 |
| Qwen-7B | 2 |
| GPT-3.5 | 1 |
Our analysis involves aggregating the scoring variances observed across all event news for the various Large Language Models (LLMs) under consideration. This comprehensive synthesis not only highlights the diversity in LLM responses but also provides a macroscopic view of their evaluative consistency and potential discrepancies. The aggregated data, which encapsulate the variance in scoring for each news event by different LLMs, are systematically presented in Table 5. This table serves as a pivotal reference point for understanding the range and distribution of LLM evaluations, offering valuable insights into their interpretative frameworks and the reliability of their analyses.
| Model | Variance |
| GLM-4 | 0.59798884 |
| ChatGLM3-6B | 0.707638507 |
| Qwen-72B | 0.784471341 |
| Qwen-7B | 0.788077699 |
| ChatGLM3-Turbo | 1.067120654 |
| Qwen-14B | 1.211226324 |
| Qwen-max | 1.363195393 |
| GPT-4 | 1.909388332 |
| InternLM2-20B | 4.893324616 |
| GPT-3.5 | 5.277003518 |
| Baichuan2-13B | 6.157081681 |
| LLaMA2-13B | 6.881628177 |
| InternLM2-7B | 7.466014898 |
| FinQwen | 9.363445905 |
| ChatGLM2-6B | 10.03005785 |
| LLaMA2-7B | 10.61274671 |
| Xuanyuan-70B | 10.83743438 |
| Xuanyuan-13B | 19.18007393 |
| Baichuan2-7B | 28.10579705 |
Upon examining the inherent biases within individual models, our analysis proceeds to consolidate the findings from each Large Language Model (LLM) to explore their collective or differential biases towards various industries. This step is crucial for understanding not only the predispositions of individual models but also for discerning any overarching trends or anomalies in their assessments of industry-related news events. By aggregating these results, we aim to delineate the extent to which these models exhibit preferential or adverse biases towards certain industry, thereby shedding light on the potential influence of these biases on the models’ analytical outputs and reliability. The synthesis of this comprehensive analysis provides a nuanced understanding of model behavior in the context of industry-specific evaluations.
In parallel with the examination of model biases, our study also delves into the temporal evolution of Large Language Models within distinct series, attributing changes to factors such as model size or software updates. To this end, we employ box line comparison charts as a visual tool to elucidate the developmental trajectories of models within the Baichuan, GLM, and LLaMA series. These charts serve to highlight variations in model performance or bias over time, providing a clear visual representation of progression or shifts in model behavior. The comparative analyses for the Baichuan series, GLM series, and LLaMA series are depicted in Figure 30, Figure 31, and Figure 32, respectively. Through these visual comparisons, we aim to offer insights into how advancements or modifications in model architecture and capabilities influence their analytical outcomes and biases.
G.2 Analysis of COT News
To delve deeper into the underlying factors contributing to potential irrationalities in Large Language Models (LLMs), our investigation extends to the analysis of reasoning outcomes facilitated by cognitive connections. This approach is predicated on the hypothesis that the manner in which LLMs forge and utilize cognitive links during the reasoning process may shed light on their logical inconsistencies or biases. For this purpose, we have meticulously selected LLMs that have demonstrated the highest, second highest, and lowest levels of performance in response to direct prompts. This selection criterion ensures a comprehensive overview, encompassing a broad spectrum of reasoning capabilities within LLMs.
The focus of this analysis is on the ’slow thinking’ aspect of model reasoning, where deliberate and methodical processing is emphasized. By examining the variance in reasoning outcomes among these models, we aim to identify patterns or anomalies that might indicate a propensity for irrational decision-making. The results of this analysis, highlighting the variance in cognitive reasoning among the selected LLMs, are systematically presented in Table 6. Through this examination, we seek to uncover the intricacies of cognitive processing in LLMs and their implications for model reliability and rationality.
| Model | Direct | COT |
| GLM-4 | 0.59798884 | 5.381799977 |
| Qwen-7B | 0.788077699 | 7.68548407 |
| ChatGLM3-Turbo | 1.067120654 | 5.670563704 |
| Xuanyuan-13B | 19.18007393 | 17.83545943 |
| Baichuan2-7B | 28.10579705 | 12.65975644 |
At the same time, we further analyzed the new5 with significant differences in ratings among different LLMs, and used the keyword detection method Bertopic to cluster and analyze the reasoning results of the models. Before clustering, the model scores and specific information of the company were removed from the reasoning results. The inference decibel of each model is clustered into 10 categories, and the distribution of scores for each category is as follows.
We further analyze the reasoning texts of the best performing GLM-4 and the worst performing Baichuan2-7B, clustering them into 10 groups with 10 keywords in each group. The key vocabulary of the two models will be summarized and a word cloud will be drawn. The results are shown in Figure 39 and Figure 39.
G.3 Analysis of Interactions
We process and analyze the information related to the interaction between the company and shareholders in a similar way.
ıin 0,2,4,…,8
G.4 Analysis of Risk-preference Questions
In the exploration of bias detection concerning risk preferences within Large Language Models (LLMs), our initial approach involved subjecting each model to three distinct input methodologies. The outcomes of these initial tests, aimed at gauging the models’ inherent risk preferences, are meticulously documented in Table 10. This foundational analysis sets the stage for more nuanced investigations into model behaviors under specific conditions.
Subsequently, our focus shifted to the models’ adherence to explicit instructions regarding risk aversion. By inputting the directive "You are a risk averse person," we were able to quantify each model’s compliance through the risk averse ratio, the details of which are encapsulated in Table 7. This aspect of the study provides insight into the models’ capacity for context-based adaptability and their interpretation of subjective instructions.
Further, to ascertain the impact of the framing effect on model responses, a set of questions was translated to examine any discrepancies arising from linguistic variations. The findings from this segment of the study, highlighting the influence of translation on model outputs, are presented in Table 8. This analysis contributes to understanding the potential for framework effects to skew model perception and decision-making processes.
Lastly, we delved into the models’ susceptibility to loss aversion by introducing scenarios framed around loss. The extent of loss aversion bias manifesting in the models’ responses was rigorously analyzed, with the summarized results being showcased in Table 9. Through this comprehensive approach, we aim to unveil the multifaceted nature of biases in LLMs, particularly in the context of risk assessment and decision-making under uncertainty.
| Model | Risk-aversion (%) |
| GPT-4 | 89.5 |
| Qwen-max | 83.5 |
| GLM-4 | 88.0 |
| Qwen-72B | 79.0 |
| ChatGLM3-Turbo | 62.5 |
| Xuanyuan-70B | 59.5 |
| Qwen-14B | 66.0 |
| InternLM2-7B | 42.5 |
| Baichuan2-13B | 44.0 |
| FinQwen | 45.0 |
| Xuanyuan-13B | 43.0 |
| ChatGLM3-6B | 53.0 |
| InternLM2-20B | 53.0 |
| Qwen-7B | 52.0 |
| Baichuan2-7B | 38.0 |
| GPT-3.5 | 37.5 |
| ChatGLM2-6B | 35.0 |
| Model | Difference (%) |
| GPT-4 | 23.5 |
| ChatGLM3-6B | 25.0 |
| Qwen-max | 28.5 |
| GLM-4 | 28.0 |
| Xuanyuan-70B | 36.0 |
| GPT-3_5 | 33.0 |
| Qwen-7B | 42.0 |
| InternLM2-7B | 43.5 |
| ChatGLM3-Turbo | 46.5 |
| FinQwen | 49.0 |
| ChatGLM2-6B | 51.0 |
| Xuanyuan-13B | 51.5 |
| Baichuan2-13B | 54.5 |
| InternLM2-20B | 48.5 |
| Qwen-72B | 56.0 |
| Baichuan2-7B | 59.0 |
| Qwen-14B | 65.0 |
| Model | Risk-aversion (%) |
| Qwen-14B | 51.0 |
| GPT-3_5 | 52.0 |
| FinQwen | 57.5 |
| Baichuan2-7B | 58.5 |
| InternLM2-7B | 60.5 |
| Qwen-max | 62.0 |
| Baichuan2-13B | 63.0 |
| InternLM2-20B | 63.0 |
| ChatGLM2-6B | 61.5 |
| Qwen-72B | 65.5 |
| Xuanyuan-13B | 66.5 |
| GLM-4 | 69.0 |
| ChatGLM3-Turbo | 74.0 |
| Xuanyuan-70B | 74.5 |
| Qwen-7B | 72.5 |
| ChatGLM3-6B | 75.0 |
| GPT-4 | 84.0 |
| Model | Input Form | Risk Averter | Risk Neutral | Risk Lover |
| Baichuan2-7B | Direct | 67 | 35 | 98 |
| Baichuan2-7B | Instruct | 76 | 24 | 100 |
| Baichuan2-7B | Translation | 67 | 58 | 75 |
| Qwen-72B | Direct | 81 | 79 | 40 |
| Qwen-72B | Instruct | 160 | 32 | 8 |
| Qwen-72B | Translation | 85 | 45 | 70 |
| Qwen-14B | Direct | 52 | 46 | 102 |
| Qwen-14B | Instruct | 134 | 27 | 39 |
| Qwen-14B | Translation | 112 | 15 | 73 |
| GLM-4 | Direct | 88 | 32 | 80 |
| GLM-4 | Instruct | 178 | 12 | 10 |
| GLM-4 | Translation | 92 | 39 | 69 |
| ChatGLM2-6B | Direct | 73 | 13 | 114 |
| ChatGLM2-6B | Instruct | 70 | 17 | 113 |
| ChatGLM2-6B | Translation | 101 | 23 | 76 |
| ChatGLM3-6B | Direct | 100 | 18 | 82 |
| ChatGLM3-6B | Instruct | 107 | 21 | 72 |
| ChatGLM3-6B | Translation | 99 | 28 | 73 |
| Xuanyuan-70B | Direct | 99 | 24 | 77 |
| Xuanyuan-70B | Instruct | 120 | 24 | 56 |
| Xuanyuan-70B | Translation | 90 | 20 | 90 |
| InternLM2-7B | Direct | 71 | 70 | 59 |
| InternLM2-7B | Instruct | 86 | 69 | 45 |
| InternLM2-7B | Translation | 81 | 63 | 56 |
| Baichuan2-13B | Direct | 76 | 16 | 108 |
| Baichuan2-13B | Instruct | 89 | 32 | 79 |
| Baichuan2-13B | Translation | 61 | 39 | 100 |
| Qwen-7B | Direct | 95 | 27 | 78 |
| Qwen-7B | Instruct | 105 | 28 | 67 |
| Qwen-7B | Translation | 106 | 24 | 70 |
| InternLM2-20B | Direct | 76 | 76 | 48 |
| InternLM2-20B | Instruct | 108 | 63 | 29 |
| InternLM2-20B | Translation | 73 | 78 | 49 |
| ChatGLM3-Turbo | Direct | 98 | 45 | 57 |
| ChatGLM3-Turbo | Instruct | 127 | 48 | 25 |
| ChatGLM3-Turbo | Translation | 62 | 85 | 53 |
| GPT-3.5 | Direct | 54 | 35 | 111 |
| GPT-3.5 | Instruct | 75 | 24 | 101 |
| GPT-3.5 | Translation | 56 | 27 | 117 |
| FinQwen | Direct | 65 | 42 | 93 |
| FinQwen | Instruct | 91 | 34 | 75 |
| FinQwen | Translation | 101 | 46 | 53 |
| GPT-4 | Direct | 118 | 41 | 41 |
| GPT-4 | Instruct | 181 | 11 | 8 |
| GPT-4 | Translation | 128 | 39 | 33 |
| Xuanyuan-13B | Direct | 83 | 26 | 91 |
| Xuanyuan-13B | Instruct | 86 | 24 | 90 |
| Xuanyuan-13B | Translation | 106 | 13 | 81 |
| Qwen-max | Direct | 74 | 89 | 37 |
| Qwen-max | Instruct | 169 | 26 | 5 |
| Qwen-max | Translation | 99 | 62 | 39 |